# An exact analytical solution for the weakly magnetized flow around an axially symmetric paraboloid, with application to magnetosphere models

### v4

@@ Appendix D The renormalized second Euler potential for the compressible case
$\hat{\beta}_{2}=$ $\ -\int_{\rho_{0}}^{\rho}\left(\frac{\big{[}a^{4}+({\rho^{\prime}}^{2}-a^{2})^{2}\,{\rho^{\prime}}^{2}\big{]}^{2}}{{\big{[}{\rho^{\prime}}^{2}+({\rho^{\prime}}^{2}-a^{2})^{2}\big{]}\,({\rho^{\prime}}^{2}-a^{2})^{2}\,\rho^{\prime}}^{3}}\right){\rm d}\rho^{\prime}=\biggl{(}\frac{1}{2\,a^{4}}-\frac{3}{a^{2}}+4\biggr{)}\,Q(\rho,a;\rho_{0})+$
@@
$L_{2}=$ $\ z-\frac{1}{2}\,\ln\biggl{(}\frac{(r+z)^{2}\,(r+z-1)}{8}\biggr{)}-\biggl{(}\frac{1}{2\,a^{4}}-\frac{3}{a^{2}}+4\biggr{)}\,Q(\rho,a;a)-\frac{1}{2\,(r+z-1)\,\rho^{2}}+$
@@
$\beta^{\mathrm{C}}_{\mathrm{ren}}=$ $\ \underbrace{(n_{\mathrm{sp}}+\nu_{1}+\nu_{2})}_{=1}\,z+\frac{n_{\mathrm{sp}}\ln(r+z-1)}{2}-\frac{\nu_{1}}{2}\,\ln\left(\frac{r+z}{2}\right)+\nu_{2}\,\Biggl{[}-\frac{1}{2}\,\ln\biggl{(}\frac{(r+z)^{2}\,(r+z-1)}{8}\biggr{)}-\Biggr{.}$

### v5

@@ Appendix D The renormalized second Euler potential for the compressible case
$\hat{\beta}_{2}=$ $\ -\int_{\rho_{0}}^{\rho}\left(\frac{\big{[}a^{4}+({\rho^{\prime}}^{2}-a^{2})^{2}\,{\rho^{\prime}}^{2}\big{]}^{2}}{{\big{[}{\rho^{\prime}}^{2}+({\rho^{\prime}}^{2}-a^{2})^{2}\big{]}\,({\rho^{\prime}}^{2}-a^{2})^{2}\,\rho^{\prime}}^{3}}\right){\rm d}\rho^{\prime}=\biggl{(}\frac{1}{2\,a^{4}}-\frac{3}{a^{2}}+4\biggr{)}\,Q(\rho,a;\rho_{0})$
@@
$L_{2}=$ $\ z-\frac{1}{2}\,\ln\biggl{(}\frac{(r+z)^{2}\,(r+z-1)}{8}\biggr{)}-\biggl{(}\frac{1}{2\,a^{4}}-\frac{3}{a^{2}}+4\biggr{)}\,Q(\rho,a;a)-\frac{1}{2\,(r+z-1)\,\rho^{2}}$
@@
$\beta^{\mathrm{C}}_{\mathrm{ren}}=$ $\ \underbrace{(n_{\mathrm{sp}}+\nu_{1}+\nu_{2})}_{=1}\,z+\frac{n_{\mathrm{sp}}\ln(r+z-1)}{2}-\frac{\nu_{1}}{2}\,\ln\left(\frac{r+z}{2}\right)+\nu_{2}\,\Biggl{[}-\frac{1}{2}\,\ln\biggl{(}\frac{(r+z)^{2}\,(r+z-1)}{8}\biggr{)}\Biggr{.}$